# Most incompatible measurements and sum-of-squares optimisation

### v1

@@ V.1 Degree four and mutually unbiased bases
to derive bounds on $\eta^{\mathrm{d}}$ valid only for MUBs. Up to degree two, everything remains unchanged with respect to Section˜III.1. For degree three, the normalisation is still given in Eq.˜32 but for the marginals, instead of the lower bound proposed in Eq.˜34, the mutual unbiasedness yields
@@
Going beyond monomials, as written in Eq.˜45, gives access to even better bounds. Numerical values for small dimensions are given in Table˜2, where we highlight some small cases for which this approach outperforms DSFB19 . In general, however, this is not the case, and understanding why in the next subsection will unlock a more general shape for the universal parent measurements that are at the centre of our work.
$\leavevmode\nobreak\ k\leavevmode\nobreak\$$\leavevmode\nobreak\ d\leavevmode\nobreak\$2 3 4 5 6
@@
V.2 Normalisable polynomials and semidefinite programming
@@
where the product is expanded from left to right. Interestingly, $\mathcal{G}_{\vec{\jmath}}^{(1)}$ contains monomials of degree up to 2k12k-1, which soon grows past the limit of degree three that was explored so far for general measurements. The difference lies in the form chosen: Eq.˜50 is not anymore a (positive) polynomial in $S_{\vec{\jmath}}$ but simply a (positive) combination of normalisable monomials in the projectors $P_{j_{x}|x}$.
@@
It is straightforward that all universal parent measurements presented so far in this work only involve normalisable polynomials. In light of the more general proof of positivity of Eq.˜50, we can now generalise the process of finding universal lower bound on $\eta^{\mathrm{g}}$ to an SOS optimisation problem TPBA24 . More precisely, for $t\geqslant 1$, denoting $\mathrm{SOS}_{2t}$ the set of sum-of-squares polynomials of degree up to $2t$ in the projectors $P_{j_{x}|x}$, we can write
@@
Eq.˜51 is simply obtained by rewriting Eq.˜13 when fixing the form of the parent measurement as a normalisable polynomial in the projectors $P_{j_{x}|x}$. Importantly, the polynomial $G$ has to be normalisable, which can be enforced by choosing a basis of normalisable monomials. In the last constraint, without loss of generality, we can set $a=1$.
Note that the similarity between Sections˜III.1 and III.3 (also Sections˜IV.1 and IV.2) generalises, so that bounds on $\chi^{\mathrm{g}}_{k}(n)$ (for $n$-outcome measurements) can be transposed to $H^{\mathrm{g}}_{k}(n)$ (for $n$-dimensional measurements) by considering rank-one measurements instead of projective ones and adding a factor $t_{\vec{\jmath}}$ in the parent measurement.
A crucial open question that numerical investigation should clarify is whether the hierarchy converges to the actual value of $\chi^{\mathrm{g}}_{k}(n)$ or $H^{\mathrm{g}}_{k}(d)$, that is, whether Eq.˜51 becomes tight for some $t$ (or even asymptotically). As shown in the literature and in this work, the answer is positive in some simple cases:
@@
V.3 Pinched-marginalisable polynomials and preliminary numerical results
As mentioned above, solving Eq.˜51 is computationally demanding. Here we add some restrictions on the form of $G$ to define a less general hierarchy for which we can derive numerical results, hence demonstrating the potential of Eq.˜51. The main idea is to select only certain types of monomials for which we can easily deal with the marginals.
@@ V.3 Pinched-marginalisable polynomials and preliminary numerical results
The restriction of this section aims at using such rewriting rules for the marginals as well, making the last constraint from Eq.˜51 much simpler as it will only depend on $k$ variable, and not all $n^{k}$ of them. More precisely, we define pinched-marginalisable monomials to be such that, for all $x$, their marginals along $x$ (when summing over all indices except for $j_{x}$) can be rewritten into a multiple of $P_{j_{x}|x}$ using the projector rules together with Eq.˜33, which we recall for convenience: $\sum_{j_{1}}P_{j_{1}|1}P_{j_{2}|2}P_{j_{1}|1}\geqslant\frac{1}{n}P_{j_{2}|2}$ (note that using Eq.˜33 repeatedly is permitted since $B\geqslant 0$ implies $ABA\geqslant 0$ for $A$ Hermitian). We denote $\Sigma_{x}$ the associated transformation. A pinched-marginalisable polynomial only has pinched-marginalisable monomials and $\Sigma_{x}$ extends to polynomials by linearity. For instance, $\Sigma_{1}(P_{j_{1}|1}P_{j_{2}|2}P_{j_{1}|1})=n^{k-2}$ and $\Sigma_{2}(P_{j_{1}|1}P_{j_{2}|2}P_{j_{1}|1})=n^{k-3}$. Here, $\Sigma_{1}$ does not use Eq.˜33 whereas $\Sigma_{2}$ does; since Eq.˜33 is an inequality, special care has to be taken to ensure the positivity of the coefficient in the polynomial. We do not explicitly state these conditions in the following but they should not be dropped in practice.
@@
For $d$-dimensional measurements, the procedure mentioned above works as well, the only difference being that Eq.˜40 is now used instead of Eq.˜33. Numerical results obtained with this hierarchy are given in Table˜1b.
VI Conclusion
@@
Cloning machine x $\frac{13}{25}$ =0.5200=0.5200 (Des22, , Eq. (11))
$t_{\vec{\jmath}}\big{(}T_{\vec{\jmath}}-\frac{1}{2}\mathds{1}\big{)}^{2}$ 2 $\frac{1}{2}$ =0.5000=0.5000 Eq.˜17
$t_{\vec{\jmath}}T_{\vec{\jmath}}\big{(}T_{\vec{\jmath}}-\frac{7-\sqrt{13}}{4}\mathds{1}\big{)}^{2}$ 3  $\frac{7+\sqrt{13}}{20}$ $\approx 0.5303$ Eq.˜30
@@
Pinched-marginalisable 4 $\approx 0.5367$ Section˜V.3
@@
Universal Normalisable ? ? Section˜V.2
 $\big{(}S_{\vec{\jmath}}-\frac{3-\sqrt{6}}{2}\mathds{1}\big{)}^{2}\big{(}S_{\vec{\jmath}}-\frac{5}{4}\mathds{1}\big{)}^{2}$ 4 $\frac{3+\sqrt{6}}{10}$ $\approx 0.5449$ Eq.˜45
$X_{\vec{\jmath}}^{\dagger}X_{\vec{\jmath}}+\text{perm.}$ 9 $\approx 0.5439$ Eq.˜50
@@
Table 3: Summary of different parent measurements for the case $d=4$ and $k=5$, both universal (for all sets of five four-dimensional measurements) and only for MUBs. When these parent measurements are polynomials in the projectors of the rank-one measurements considered, we indicate the corresponding degree. All values are analytical, except for the pinched-marginalisable ones, which we obtain by solving the corresponding SDP (52) numerically. Higher values could be reached (say, with degree 8) but this was out of reach with our code; also, we did not implement the generic hierarchy for normalisable parent measurements. Apart from the cloning machine, all universal bounds have their equivalent for four-outcome measurements, dropping the factor $t_{\vec{\jmath}}$ and replacing $T_{\vec{\jmath}}$ by $S_{\vec{\jmath}}$. For MUBs, the value with degree 2k1=92k-1=9 is obtained by sums of permutations of Eq.˜50 as described in (DSFB19, , Eq. (13)); the fact that it is lower than the value reached above with degree four is not true in general, see Table˜2. In the last line, we give the exact value for MUBs; $\Pi_{\vec{\jmath}}$ is the eigenprojector of the maximal eigenvalue of $S_{\vec{\jmath}}$, and it is shown in (DSFB19, , Appendix D.1) that this parent measurement is optimal and asymptotically reached by powers of $S_{\vec{\jmath}}$.
@@ ACKNOWLEDGEMENTS.
We thank Armin Tavakoli for suggesting to study Pauli strings, which was the starting point of this work, Gabriel Cobucci for sharing related notes, Mateus Araújo for helping out using Moment at an early stage of the numerical experiments, and Máté Farkas for useful discussions. Research reported in this paper was supported through the Research Campus Modal funded by the German Federal Ministry of Education and Research (fund numbers 05M14ZAM, 05M20ZBM), and partly funded within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733 (VERIqTAS).

### v2

@@ VI.1 Degree four and mutually unbiased bases
to derive bounds on $\eta^{\mathrm{d}}$ valid only for MUBs. Up to degree two, everything remains unchanged with respect to Section˜IV.1. For degree three, the normalisation is still given in Eq.˜40 but for the marginals, instead of the lower bound proposed in Eq.˜42, the mutual unbiasedness yields
@@
Going beyond monomials, as written in Eq.˜53, gives access to even better bounds. Numerical values for small dimensions are given in Table˜2, where we highlight some small cases for which this approach outperforms DSFB19 . In general, however, this is not the case, and understanding why in the next subsection will unlock a more general shape for the universal parent measurements that are at the centre of our work.
$\penalty 10000\ k\penalty 10000\$$\penalty 10000\ d\penalty 10000\$2 3 4 5 6
@@
VI.2 Normalisable polynomials and semidefinite programming
@@
where the product is expanded from left to right. Interestingly, $\mathcal{G}_{\vec{\jmath}}^{(1)}$ contains monomials of degree up to 2k12k-1, which soon grows past the limit of degree three that was explored so far for general measurements. The difference lies in the form chosen: Eq.˜58 is not anymore a (positive) polynomial in $S_{\vec{\jmath}}$ but simply a (positive) combination of normalisable monomials in the projectors $P_{j_{x}|x}$.
@@
It is straightforward that all universal parent measurements presented so far in this work only involve normalisable polynomials. In light of the more general proof of positivity of Eq.˜58, we can now generalise the process of finding universal lower bound on $\eta^{\mathrm{g}}$ to an SOS optimisation problem TPBA24 . More precisely, for $t\geqslant 1$, denoting $\mathrm{SOS}_{2t}$ the set of sum-of-squares polynomials of degree up to $2t$ in the projectors $P_{j_{x}|x}$, we can write
@@
Eq.˜59 is simply obtained by rewriting Eq.˜13 when fixing the form of the parent measurement as a normalisable polynomial in the projectors $P_{j_{x}|x}$. Importantly, the polynomial $G$ has to be normalisable, which can be enforced by choosing a basis of normalisable monomials. In the last constraint, without loss of generality, we can set $a=1$.
Note that the similarity between Sections˜IV.1 and IV.3 (also Sections˜V.1 and V.2) generalises, so that bounds on $\chi^{\mathrm{g}}_{k}(n)$ (for $n$-outcome measurements) can be transposed to $H^{\mathrm{g}}_{k}(n)$ (for $n$-dimensional measurements) by considering rank-one measurements instead of projective ones and adding a factor $t_{\vec{\jmath}}$ in the parent measurement.
A crucial open question that numerical investigation should clarify is whether the hierarchy converges to the actual value of $\chi^{\mathrm{g}}_{k}(n)$ or $H^{\mathrm{g}}_{k}(d)$, that is, whether Eq.˜59 becomes tight for some $t$ (or even asymptotically). As shown in the literature and in this work, the answer is positive in some simple cases:
@@
VI.3 Pinched-marginalisable polynomials and preliminary numerical results
As mentioned above, solving Eq.˜59 is computationally demanding. Here we add some restrictions on the form of $G$ to define a less general hierarchy for which we can derive numerical results, hence demonstrating the potential of Eq.˜59. The main idea is to select only certain types of monomials for which we can easily deal with the marginals.
@@ VI.3 Pinched-marginalisable polynomials and preliminary numerical results
The restriction of this section aims at using such rewriting rules for the marginals as well, making the last constraint from Eq.˜59 much simpler as it will only depend on $k$ variable, and not all $n^{k}$ of them. More precisely, we define pinched-marginalisable monomials to be such that, for all $x$, their marginals along $x$ (when summing over all indices except for $j_{x}$) can be rewritten into a multiple of $P_{j_{x}|x}$ using the projector rules together with Eq.˜41, which we recall for convenience: $\sum_{j_{1}}P_{j_{1}|1}P_{j_{2}|2}P_{j_{1}|1}\geqslant\frac{1}{n}P_{j_{2}|2}$ (note that using Eq.˜41 repeatedly is permitted since $B\geqslant 0$ implies $ABA\geqslant 0$ for $A$ Hermitian). We denote $\Sigma_{x}$ the associated transformation. A pinched-marginalisable polynomial only has pinched-marginalisable monomials and $\Sigma_{x}$ extends to polynomials by linearity. For instance, $\Sigma_{1}(P_{j_{1}|1}P_{j_{2}|2}P_{j_{1}|1})=n^{k-2}$ and $\Sigma_{2}(P_{j_{1}|1}P_{j_{2}|2}P_{j_{1}|1})=n^{k-3}$. Here, $\Sigma_{1}$ does not use Eq.˜41 whereas $\Sigma_{2}$ does; since Eq.˜41 is an inequality, special care has to be taken to ensure the positivity of the coefficient in the polynomial. We do not explicitly state these conditions in the following but they should not be dropped in practice.
@@
For $d$-dimensional measurements, the procedure mentioned above works as well, the only difference being that Eq.˜48 is now used instead of Eq.˜41. Numerical results obtained with this hierarchy are given in Table˜1b.
VII Conclusion
@@
Cloning machine x $\frac{13}{25}$ =0.5200=0.5200 Eq.˜17
$t_{\vec{\jmath}}\big(T_{\vec{\jmath}}-\frac{1}{2}\mathds{1}\big)^{2}$ 2 $\frac{1}{2}$ =0.5000=0.5000 Eq.˜25
$t_{\vec{\jmath}}T_{\vec{\jmath}}\big(T_{\vec{\jmath}}-\frac{7-\sqrt{13}}{4}\mathds{1}\big)^{2}$ 3 $\frac{7+\sqrt{13}}{20}$ $\approx 0.5303$ Eq.˜38
@@
Pinched-marginalisable 4 $\approx 0.5367$ Section˜VI.3
@@
Universal Normalisable ? ? Section˜VI.2
$\big(S_{\vec{\jmath}}-\frac{3-\sqrt{6}}{2}\mathds{1}\big)^{2}\big(S_{\vec{\jmath}}-\frac{5}{4}\mathds{1}\big)^{2}$ 4 $\frac{3+\sqrt{6}}{10}$ $\approx 0.5449$ Eq.˜53
$X_{\vec{\jmath}}^{\dagger}X_{\vec{\jmath}}+\text{perm.}$ 9 $\approx 0.5439$ Eq.˜58
@@
Table 3: Summary of different parent measurements for the case $d=4$ and $k=5$, both universal (for all sets of five four-dimensional measurements) and only for MUBs. When these parent measurements are polynomials in the projectors of the rank-one measurements considered, we indicate the corresponding degree. All values are analytical, except for the pinched-marginalisable ones, which we obtain by solving the corresponding SDP (60) numerically. Higher values could be reached (say, with degree 8) but this was out of reach with our code; also, we did not implement the generic hierarchy for normalisable parent measurements. Apart from the cloning machine, all universal bounds have their equivalent for four-outcome measurements, dropping the factor $t_{\vec{\jmath}}$ and replacing $T_{\vec{\jmath}}$ by $S_{\vec{\jmath}}$. For MUBs, the value with degree 2k1=92k-1=9 is obtained by sums of permutations of Eq.˜58 as described in (DSFB19, , Eq. (13)); the fact that it is lower than the value reached above with degree four is not true in general, see Table˜2. In the last line, we give the exact value for MUBs; $\Pi_{\vec{\jmath}}$ is the eigenprojector of the maximal eigenvalue of $S_{\vec{\jmath}}$, and it is shown in (DSFB19, , Appendix D.1) that this parent measurement is optimal and asymptotically reached by powers of $S_{\vec{\jmath}}$.
@@ ACKNOWLEDGEMENTS.
We thank Armin Tavakoli for suggesting to study Pauli strings, which was the starting point of this work, Gabriel Cobucci for sharing related notes, Mateus Araújo for helping out using Moment at an early stage of the numerical experiments, Andreas Bluhm for pointing out relevant references, and Nicola d’Alessandro and Máté Farkas for useful discussions. Research reported in this paper was supported through the Research Campus Modal funded by the German Federal Ministry of Education and Research (fund numbers 05M14ZAM, 05M20ZBM), and partly funded within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733 (VERIqTAS).